OpenStudy (anonymous):

y + 12 = x^2 + x What are the x-intercepts

OpenStudy (anonymous):

make y=0, and solve for x

jhonyy9 (jhonyy9):

(x-3)(x+4)=0

OpenStudy (anonymous):

Could you explain how do you factor out that expression jhonny9?

jhonyy9 (jhonyy9):

calcule these you to prove or check these

jhonyy9 (jhonyy9):

so or you dont believe me ?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

I don't how to factor.

jhonyy9 (jhonyy9):

and how you see is it right ?

OpenStudy (anonymous):

I used the quadractic formula instead

jhonyy9 (jhonyy9):

No-data attention so because there will be x2 +x -12 =0

jhonyy9 (jhonyy9):

so than you need calcule the +1x from two numbers what if you assum will add to you +1x and two numbers what if you multiply will get -12

OpenStudy (anonymous):

ahh -3 and 4

jhonyy9 (jhonyy9):

so after you will check or will make the calcule will get that 4-3=+1 and 4*(-3)=-12

jhonyy9 (jhonyy9):

right ?

OpenStudy (anonymous):

and then I form (x-4)(x+3)

OpenStudy (anonymous):

Yeah, right!

OpenStudy (anonymous):

Thank U =)

jhonyy9 (jhonyy9):

yes hope so much that was understandably sure

jhonyy9 (jhonyy9):

yw